what is 26 1/6 minus 6 13/16
Found 2 solutions by rfer, shree840:
Answer by rfer(16322)

26 1/6 - 6 13/16
26 8/48 - 6 39/48
25 56/48 - 6 39/48 = 19 17/48

Answer by shree840(260)

26 1/6
- 6 13/16
lcm of 6 and 16 is 3*2 2*2*2*2
we have one 2 common
for lcm we take one of common and take everything else
lcm =2*3*2*2*2=48
rewrite original expression 26 8/48
- 6 39/48
since 8 is less than 39 we borrow 1 which is 48/48
top one is 25 56/48
bottom 6 39/48
(subtract) 19 17/48